Simplify:
x = 7, y = -2, and z = -4
-x + y + z

Answer by stanbon(75887)

Simplify:
x = 7, y = -2, and z = -4
-x + y + z = -(7) + (-2) + (-4) = -7-2-4 = -13
-----------------
Cheers,
Stan H.
------------